a train moving away from a sound detector at 30.0 m/s blows a 300-hz whistle.speed of sound is 340 m/s.what frequency is detected on each of the following?

a.stationary train.

b.a train moving away from the first train at speed of 20.0 m/


1
Expert's answer
2021-11-22T10:12:20-0500

(a) We can find the frequency detected from the Doppler Shift formula:


"f_o=f_s(\dfrac{v}{v+v_s}),""f_o=300\ Hz\times(\dfrac{340\ \dfrac{m}{s}}{340\ \dfrac{m}{s}+30\ \dfrac{m}{s}})=276\ Hz."

(b) We can find the frequency detected from the Doppler Shift formula:


"f_o=f_s(\dfrac{v-v_o}{v+v_s}),""f_o=300\ Hz\times(\dfrac{340\ \dfrac{m}{s}-20\ \dfrac{m}{s}}{340\ \dfrac{m}{s}+30\ \dfrac{m}{s}})=259.5\ Hz."